What is the value of k if one root root of (3k+1)x^2=-5 is 1? Please help me :)
Answer by stanbon(75887)

What is the value of k if one root root of (3k+1)x^2=-5 is 1? Please help me :)
------------
Rearrange:
f(x) = (3k+1)x^2+5
If 1 is a root, f(1) = zero
But f(1) = (3k+1)*1^2+5 = 0
3k+1+5 = 0
3k = -6
k = -2
==============
Cheers,
Stan H.